A 〖MgCO〗_3 solution was prepared by dissolving 0.500 g 〖MgCO〗_3 in sufficient water to produce 250.00 mL solution. A 20.00 mL aliquot is taken and put into an Erlenmeyer flask. A 25.00 mL of 0.1500 M 〖HNO〗_3 is added into the flask. The resulting solution is then titrated with 0.12 M NaOH. Calculate the volume of NaOH required for the back titration.
1
Expert's answer
2020-12-10T07:48:19-0500

1) MgCO3 + 2HNO3 = MgNO3 + H2CO3

2) HNO3 + NaOH = NaNO3 + H2O


M (MgCO3) = 84.31 g/mol

n (MgCO3) = 0.500 / 84.31 = 0.006 mol

CM (MgCO3) = n/V = 0.006 / 0.25 = 0.024 M


20.00 ml of this MgCO3 solution contains:

n1(MgCO3) = 20.00/1000 x 0.024 = 0.0005 mol

According to the equation 1, n(HNO3) = 2 x n(MgCO3) = 2 x 0.0005 = 0.001 mol


n(HNO3)actual = 0.1500 x 25.00/1000 = 0.0038 mol

So that present HNO3 is in excess: 0.0038 - 0.001 = 0.0028 mol


According to the eqanion 2, n (HNO3) = n (NaOH) = 0.0028 mol

CM=n/V

V (NaOH) = n/CM = 0.0028 / 0.12 = 0.03 L = 30 ml